# Not All Frame Features Are Equal: Video-to-4D Generation via Decoupling Dynamic-Static Features

### v1

@@ Abstract
Recently, the generation of dynamic 3D objects from a video has shown impressive results. Existing methods directly optimize Gaussians using whole information in frames. However, when dynamic regions are interwoven with static regions within frames, particularly if the static regions account for a large proportion, existing methods often overlook information in dynamic regions and are prone to overfitting on static regions. This leads to producing results with blurry textures. We consider that decoupling dynamic-static features to enhance dynamic representations can alleviate this issue. Thus, we propose a dynamic-static feature decoupling module (DSFD). Along temporal axes, it regards the portions of current frame features that possess significant differences relative to reference frame features as dynamic features. Conversely, the remaining parts are the static features. Then, we acquire decoupled features driven by dynamic features and current frame features. Moreover, to further enhance the dynamic representation of decoupled features from different viewpoints and ensure accurate motion prediction, we design a temporal-spatial similarity fusion module (TSSF). Along spatial axes, it adaptively selects a similar information of dynamic regions. Hinging on the above, we construct a novel approach, DS4D. Experimental results verify our method achieves state-of-the-art (SOTA) results in video-to-4D. In addition, the experiments on a real-world scenario dataset demonstrate its effectiveness on the 4D scene. Our code will be publicly available.
@@ 1 Introduction
There are two main streams in current approaches to improving generation quality. Inference-based methods [10, 11, 12, 13] can generate high-quality 4D content by capturing temporal-spatial correlations in its 4D diffusion model. Another stream is optimization-based methods [14, 1, 15, 16]. These methods generate 4D content through distilling spatial prior knowledge from pre-trained multi-view [17, 18] diffusion models. However, the above methods only model temporal-spatial correlations using whole information in frames, but fail to explicitly differentiate between the dynamic and static regions of the frame. If static regions account for a significant portion, these models overlook dynamic information. Thus they tend to overfit static regions, resulting in a diminished capacity to perceive texture variations in dynamic zones. As shown in Fig.1 (a), the previous work (STAG4D [1]) produces results with blurred texture details (e.g., wrinkles in clothes), particularly when confronted with B-type input video.
To tackle above problems, we present a novel approach DS4D, which decouples dynamic and static features along temporal and spatial axes to enhance dynamic representations for high-quality 4D generation. In particular, we propose dynamic-static feature decoupling module (DSFD) to obtain decoupled features. We assume some portions of current frame features exhibit significant differences relative to reference frame features. Such portions always contain important knowledge of texture variations and motion trends under current timestamps. Therefore, it can be regarded as the dynamic feature of the current frame. In contrast, the remaining parts are the static features. Based on this assumption, we decompose the dynamic components between each frame feature and reference frame features along the temporal axes. The dynamic component (also termed dynamic features), as difference features between two frame features, is able to represent the dynamic information (as shown in Fig.1 (b)). At last, we acquire decoupled features driven by dynamic components and corresponding frame features.
Note that spatial occlusion results in the inability of dynamic components captured from a specific viewpoint to inadequately represent the authentic dynamic information in 4D space. To mitigate this issue, we design temporal-spatial similarity fusion module (TSSF) which engages in enhancing dynamic representations in features. The Gaussian points are initialized by utilizing a large reconstruction model [19]. Then, point features are obtained by retrieving decoupled features for Gaussian points via view projection in TSSF. Subsequently, TSSF produces fused Gaussian features by adaptively selecting information on dynamic regions containing similar texture, shape, and motion representations from point feature space at same timestamps along spatial axes. Finally, fused Gaussian features with strong dynamic representation, are used for 4D generation.
@@
  • We propose a novel framework DS4D, which decouples dynamic-static information along temporal-spatial axes to enhance dynamic representations for high-quality 4D content generation.
@@
  • Experimental results on Consistent4D dataset [20] and Objaverse dataset [21] demonstrate our DS4D outperforms other SOTA methods in terms of video quality, motion fidelity, and temporal-spatial consistency. Furthermore, the experiments on a real-world scenario dataset [2] verify its effectiveness on 4D scenes.
@@ 2 Related Works
Figure 2: Overview of our proposed DS4D. Given an input video and corresponding multi-view sequences, our DS4D decouples features of the frame at time $t$ based on the reference frame in DSFD module. Next, we acquire point features by retrieving each decoupled feature for Gaussian points via view projection, and we obtain fused Gaussian features by adaptively selecting similar dynamic information from point feature space in TSSF module. Finally, through Deformation MLP, our method generates 4D content. During the process, we utilize DINOv2 [22] to extract each frame feature and use a large reconstruction model (LRM) [19] to generate the corresponding mesh. We perform uniform sampling on the mesh to obtain point clouds. The point clouds are the initialization of static 3D Gaussian points.
@@ 3.2 Initialization
Frame Feature Extraction. The frame features extracted by a visual foundation model must incorporate structural information, thereby ensuring that the information contained within the corresponding features is utilized to decouple the dynamic and static features. Recent research [49] demonstrates DINOv2 [22] has a stronger ability to capture the structure of input image than other foundation models [50, 51]. Therefore, we leverage DINOv2 to extract frame features $\textbf{F}=\{f^{(0,0)},f^{(0,1)},\dots,f^{(i,j)},\dots,f^{(t,v)}\}$ from sequences $\textbf{I}=\{I^{(0,0)},I^{(0,1)},\dots,I^{(i,j)},\dots,I^{(t,v)}\}$, where $f^{(i,j)}\in\mathbb{R}^{P\times D}$, $P$ and $D$ are the number of tokens and token dimension, respectively. Each feature $f^{(i,j)}$ encodes the geometry and texture of the corresponding frame.
Gaussian Points Initialization. Previous works usually randomly initialize Gaussian points, resulting in unstable topology during optimization. It ultimately affects the quality of the results. Considering the above problem, we generate the point clouds from the frame at middle timestamp by using a large reconstruction model [19], which is the initialization of static 3D Gaussian points. The initialization approach provides geometric prior and ensures the stability of subsequent optimization.
@@ 3.3 Dynamic-Static Feature Decoupling (DSFD)
Now, we obtain frame features F which include the features of all frames and their corresponding multi-view features. Then, we start from the following two aspects to decouple dynamic-static features: 1) Which frame features are regarded as references; 2) How to decouple frame features along temporal axes.
Which frame features are regarded as references? In order to support the model stably predict 4D contents, the decoupled dynamic information should precisely represent the variations in motion of the current frame relative to the entire video stream. Therefore, it is essential to define the reference frames employed to decouple the motion deformation between the current frame and reference frames. The most direct way is to treat each frame feature as a reference and decouple the dynamic and static features between references and other frame features one by one under different viewpoints along the temporal axis. However, it usually consumes considerable computation time.
To alleviate this issue, we sample reference frame features from F and decouple features based on the current frame and reference frame features. This manner greatly reduces the computation time. Concretely, the frame $f^{(\frac{t}{2},j)}$ of middle timestamp and average frame features $\bar{f}^{(\bar{t},j)}$ are regarded as references. For frame $f^{(i,j)}$ at timestamp $i$, under viewpoint $j$, we can calculate the difference information of $f^{(i,j)}$ relative to $f^{(\frac{t}{2},j)}$ and $\bar{f}^{(\bar{t},j)}$. Based on the difference information, we decouple the dynamic and static features.
@@
How to decouple frame features? To obtain decoupled features, we propose the dynamic-static feature decoupling module (DSFD). In DSFD, we aim to calculate differences by effectively utilizing both geometric and texture information present in features while maintaining the original distribution of pre-trained DINOv2 model. To achieve this, we implement a parameter-free architecture for decoupling.
The parameter-free decoupling architecture is shown in Fig.3. Specifically, given specific $j$-th view in timestamp $i$, we perform the projection of frame features $f^{(i,j)}$ onto reference frame features $r^{j}$. Actually, the projection serves as the component with similar information between the two frame features. The similar feature zones belong to the regions where there are no texture and shape variations between the two frames. Thus, the component can represent static features. Formally, for each frame feature $f^{(i,j)}$, the static features become:
$f_{\text{static}}^{(i,j)}=f^{(i,j)}\cdot r^{j}\frac{r^{j}}{\|r^{j}\|_{2}^{2}},$ (1)
where $f^{(i,j)}\cdot r^{j}=|f^{(i,j)}|\times|r^{j}|\times\cos(\theta)$, and $\theta$ is included angle between $f^{(i,j)}$ and $r^{j}$. Through projection, we obtain static features that exhibit the same geometric and texture information between current frame and reference frames.
We calculate the differential between current frame features and static features to achieve dynamic features (also termed dynamic components):
$f_{\text{dynamic}}^{(i,j)}=f^{(i,j)}-f_{\text{static}}^{(i,j)}.$ (2)
The dynamic features possess significant difference information, which represent the motion knowledge of the current frame relative to the reference frame. Therefore, we leverage dynamic features to enhance the dynamic representations in the current frame. In particular, for each type of reference frame features $r^{j}$, including $f^{(\frac{t}{2},j)}$ and $\bar{f}^{(\bar{t},j)}$, we perform eq. 1, eq. 2 to obtain corresponding dynamic features $f_{\text{dynamic}}^{(i,j)}$ at each timestamp. Afterwards, we append to each token of current frame features $f^{(i,j)}$ with dynamic features $f_{\text{dynamic}}^{(i,j)}$. The appended features are $\alpha_{t/2}^{(i,j)}$ and $\alpha_{avg}^{(i,j)}$. Last, we calculate the average pooling between the two features and acquire features $f_{d}^{(i,j)}$ with rich dynamic representations. We integrate all the features obtained based on both temporal and spatial axes into decoupled features $\textbf{F}_{\textbf{D}}=\{f_{d}^{(0,0)},f_{d}^{(0,0)},\dots,f_{d}^{(i,j)},\dots,f_{d}^{(t,v)}\}$.
3.4 Temporal-Spatial Similarity Fusion (TSSF)
After we get decoupled features $\textbf{F}_{\textbf{D}}$, we have to enhance dynamic representations further. Because the dynamic features obtained from a particular viewpoint do not adequately represent the complete dynamic information in the 4D space. Thus, we design temporal-spatial similarity fusion module (TSSF) to select similar dynamic information from decoupled features at the same timestamp.
Inspired by [52], we retrieve decoupled features from Gaussian points by view projection. In particular, decoupled features are integrated into point features based on the given camera pose. The Gaussian points are well-aligned with decoupled features thanks to the projection operations.
Afterwards, we extract point features $\textbf{F}_{\textbf{p}}=\{f_{p}^{(0,0)},f_{p}^{(0,1)},\dots,f_{p}^{(i,j)},\dots,f_{p}^{(t,v)}\}$ which encapsulate extensive dynamic information. Among different viewpoints, we consider the texture, shape, and motion representations from point features to be similar for the same spatial area of the object in the frame. Hence, utilizing the similarities in information across viewpoints and aggregating them contributes to the enhancement of dynamic representations. According to the assumption, we propose an adaptive fusion to select such information.
@@ 3.4 Temporal-Spatial Similarity Fusion (TSSF)
Fig.4 shows the framework of adaptive fusion in TSSF. Along spatial axes, we use adaptive fusion to merge similar dynamic information from point features belonging to the same time. However, in some situations, pseudo multi-view sequences are generated by Diffusion model with real single-view video as a condition. The quality of pseudo multi-view sequences influences fusion effectiveness. For example, dynamic regions are noticeably obscured in specific viewpoints. Hence, we design two adaptive fusion methods, global awareness fusion (GA) and distance awareness fusion (DA), to cope with different scenarios.
Global awareness (GA). At time $i$, the fully-connection layer (FC) generates a score map $\textbf{W}=\{w^{(0,0)},\dots,w^{(t,v)}\}$ for the point features of each view,
@@
The zones that exhibit similar dynamic information in point features of different viewpoints hold significant weights. Hence, similar dynamic features can be merged.
Distance awareness (DA). In general, the frames of the front view are from real single-view video, and the other views are pseudo. The front frame reveals the relatively intact motion areas of the object at various timestamps compared to the other views. Thus, at time $i$, we calculate the $L_{1}$ distance between point features from the front view and point features from the other views. Then, we append pseudo point features with distance features and merge them by eq. 3 and eq. 4. The fused pseudo point features are appended to real point features. And we once again merge the features by eq. 3 and eq. 4 to acquire the fused point features $f_{a}^{i}$. Through the above pre-processing feature filtering, we reserve the distance features of the pseudo point features relative to the real point features in advance instead of complete pseudo point features. Thus, we reduce the impact of pseudo frame quality on merging dynamic information.
In addition to fused features we obtained through adaptive fusion, we also present dynamic Gaussian features as 3D Gaussian features with HexPlane [53] for inherently regularizing features of fields and guaranteeing their smoothness. Formally, the dynamic Gaussian features at time $i$:
@@
where $X_{i}=(x_{i},y_{i},z_{i})$ denotes the position of Gaussian points, $s_{i}$ and $\gamma_{i}$ denote scale and rotation at time $i$. $\sigma$ and $\zeta$ represent opacity and spherical harmonic coefficients of the radiance. Then, we combine dynamic Gaussian features $f_{\text{hg}}^{i}$ and fused point features $f_{a}^{i}$ at time $i$, and map them to fused Gaussian features via learnable linear transformation. The fused Gaussian features not only represent Gaussian intrinsic properties but also capture rich dynamic information, essential for 4D generation by accurately predicting deformation using Deformation MLP.

### v2

@@ Abstract
Recently, the generation of dynamic 3D objects from a video has shown impressive results. Existing methods directly optimize Gaussians using whole information in frames. However, when dynamic regions are interwoven with static regions within frames, particularly if the static regions account for a large proportion, existing methods often overlook information in dynamic regions and are prone to overfitting on static regions. This leads to producing results with blurry textures. We consider that decoupling dynamic-static features to enhance dynamic representations can alleviate this issue. Thus, we propose a dynamic-static feature decoupling module (DSFD). Along temporal axes, it regards the regions of current frame features that possess significant differences relative to reference frame features as dynamic features. Conversely, the remaining parts are the static features. Then, we acquire decoupled features driven by dynamic features and current frame features. Moreover, to further enhance the dynamic representation of decoupled features from different viewpoints and ensure accurate motion prediction, we design a temporal-spatial similarity fusion module (TSSF). Along spatial axes, it adaptively selects similar information of dynamic regions. Hinging on the above, we construct a novel approach, DS4D. Experimental results verify our method achieves state-of-the-art (SOTA) results in video-to-4D. In addition, the experiments on a real-world scenario dataset demonstrate its effectiveness on the 4D scene. Our code will be publicly available.
@@ 1 Introduction
There are two main streams in current approaches to improving generation quality. Inference-based methods [10, 11, 12, 13] can generate high-quality 4D content by capturing temporal-spatial correlations in its 4D diffusion model. Another stream is optimization-based methods [14, 1, 15, 16]. These methods generate 4D content through distilling spatial prior knowledge from pre-trained multi-view [17, 18] diffusion models. However, previous methods only model temporal-spatial correlations using whole information in frames, but fail to explicitly differentiate between the dynamic and static regions within a frame. If static regions account for a significant portion, previous methods overlook dynamic information. Thus they tend to overfit static regions, resulting in a diminished capacity to perceive texture variations in dynamic regions. As shown in Fig.1 (a), previous work (STAG4D [1]) produces results with blurry texture (e.g., wrinkles in clothes), particularly when confronted with B-type input video.
To tackle above problems, we present a novel approach DS4D, which decouples dynamic and static features along temporal and spatial axes to enhance dynamic representations for high-quality 4D generation. In particular, we propose dynamic-static feature decoupling module (DSFD) to obtain decoupled features. We assume some regions of current frame features exhibit significant differences relative to reference frame features. Such regions always contain important knowledge of texture, shape variations, and motion trends under current timestamps. Therefore, it can be regarded as the dynamic feature of the current frame. In contrast, the remaining parts are the static features. Based on this assumption, we decompose the dynamic components between each frame feature and reference frame features along the temporal axes. The dynamic component (also termed dynamic features), as difference features between two frame features, is able to represent the dynamic information (as shown in Fig.1 (b)). At last, we acquire decoupled features driven by dynamic components and corresponding frame features.
Note that spatial occlusion leads to the failure of dynamic components captured from a specific viewpoint to inadequately represent the inherent dynamic information in 4D space. To mitigate this issue, we design temporal-spatial similarity fusion module (TSSF) which engages in enhancing features’ dynamic representations. Firstly, Gaussian points are initialized by utilizing a large reconstruction model [19]. Then, point features are obtained by retrieving decoupled features for Gaussian points via view projection in TSSF. Subsequently, TSSF produces fused Gaussian features by adaptively selecting information on dynamic regions containing similar texture, shape, and motion representations from point feature space at same timestamps along spatial axes. Finally, fused Gaussian features with strong dynamic representation, are used for 4D generation.
@@
  • To our best knowledge, we are the first to propose a framework DS4D that decouples dynamic-static information in frames along temporal-spatial axes to enhance dynamic representations for high-quality 4D generation.
@@
  • Experimental results on Consistent4D dataset [20] and Objaverse dataset [21] demonstrate our DS4D outperforms other SOTA methods in terms of video quality, motion fidelity, and temporal-spatial consistency. Furthermore, the experiments on a complex real-world scenario dataset [2] verify its effectiveness on 4D scenes.
@@ 2 Related Works
Figure 2: Overview of our proposed DS4D. Given an input video and corresponding multi-view sequences, our DS4D decouples features of the frame at time $t$ based on the reference frame in DSFD module. Next, we acquire point features by retrieving each decoupled feature for Gaussian points via view projection, and we obtain fused Gaussian features by adaptively selecting similar dynamic information from point feature space in TSSF module. Finally, through Deformation MLP, our method generates 4D content.
@@ 3.2 Initialization
Frame Feature Extraction. Owing to the superior feature extraction capability of the visual foundation model, we employ it as the feature extractor to obtain high-quality features. Thanks to the strong feature extraction ability of DINOv2 [48], we leverage DINOv2 to extract frame features $\textbf{F}=\{f^{(0,0)},f^{(0,1)},\dots,f^{(i,j)},\dots,f^{(t,v)}\}$ from sequences $\textbf{I}=\{I^{(0,0)},I^{(0,1)},\dots,I^{(i,j)},\dots,I^{(t,v)}\}$, where $f^{(i,j)}\in\mathbb{R}^{P\times D}$, $P$ and $D$ are the number of tokens and token dimension, respectively. Each feature $f^{(i,j)}$ encodes the geometry and texture of the corresponding frame.
22footnotetext: The discussion about the impact of multi-view sequence quality and the robustness of our methods can be seen in supplementary material (Sec. 12.1).
Gaussian Points Initialization. Previous works [1, 16] initialize Gaussian points randomly, leading to unstable topology during optimization and ultimately compromising the quality of results. To solve the above problem, we initialize static 3D Gaussian points through pre-defined point clouds. Specifically, we employ a large reconstruction model [19] to generate the point clouds from the middle frame. This initialization approach provides a geometric prior and ensures the stability of subsequent optimization.
@@ 3.3 Dynamic-Static Feature Decoupling (DSFD)
We employ a set of frame features F, which consists of all frame features and their corresponding multi-view features, to decouple dynamic-static features. Specifically, our decoupling process involves two aspects: 1) Finding out reference frame features; 2) Decoupling dynamic and static features along temporal axes.
Finding out reference frame features. Decoupling frame-by-frame features usually consumes considerable computation time. Hence, we specify special frames as references for decoupling. It is critical to determine whether reference frame features can accurately represent the semantic content of the entire video. Considering that we initialize static Gaussian points at the middle frame, the whole model needs to capture the texture, shape, and motion transformation from the middle frame to any frame during training and inference. Therefore, we choose the frame features $f^{(\frac{t}{2},j)}$ of the middle timestamp as one of the reference frame features, which represent the semantic. Furthermore, to represent the motion trends in video, we need a kind of feature that can describe the motion variations of the entire video. Hence, we also specify the average frame features $\bar{f}^{(\bar{t},j)}$ as reference frame features, which represent the average motion variations.
Decoupling dynamic and static features along temporal axes. To reliably decouple dynamic and static information in frame features, we need to obtain the regions of current frame features that exhibit significant differences relative to reference frame features. These regions contain important knowledge of texture, shape variations, and motion trends under current timestamps. The regions with significant differences mean dynamic parts between the two frame features, while regions with similar information mean static parts between the two frame features. Therefore, we propose the dynamic-static feature decoupling module (DSFD), as shown in Fig.13.
Concretely, given specific $j$-th view in timestamp $i$, we project frame features $f^{(i,j)}$ onto reference frame features $r^{j}$. The projection represents the semantic overlapping part of the two frames. In other words, regions in the two frames with no texture and shape transformations are identified as the static part. The process is mathematically formulated as follows:
$f_{\text{static}}^{(i,j)}=(\frac{f^{(i,j)}\cdot r^{j}}{\|r^{j}\|_{2}})\frac{r^{j}}{\|r^{j}\|_{2}},$ (1)
where $\cdot$ denotes dot product.
The difference between current frame features and static features means the orthogonal vector. Then, we identify the orthogonal vector as the semantic difference between the current frame and the reference frame. Hence, the orthogonal vector can be regarded as dynamic features:
$f_{\text{dynamic}}^{(i,j)}=f^{(i,j)}-f_{\text{static}}^{(i,j)}.$ (2)
At each timestamp, we obtain the dynamic features in a particular view. Dynamic features contain significant difference information, which represents the motion knowledge of the current frame relative to the reference frame. Meanwhile, frame features contain texture and shape attributes of the current frame, which are useful for subsequently texture and shape learning. Hence, we leverage dynamic features to supplement dynamic information in the frame features while retaining the inherent attributes of the frame features.
Here, we append each token of the current frame features $f^{(i,j)}$ with dynamic features $f_{\text{dynamic}}^{(i,j)}$. Then, the appended features $f_{d}^{(i,j)}$ are integrated based on both the temporal $i$ and spatial $j$, which are named decoupled features $\textbf{F}_{\textbf{D}}=\{f_{d}^{(0,0)},f_{d}^{(0,0)},\dots,f_{d}^{(i,j)},\dots,f_{d}^{(t,v)}\}$, where $t$ and $v$ denote the number of frames and number of views, respectively.
3.4 Temporal-Spatial Similarity Fusion (TSSF)
Since the features in $\textbf{F}_{\textbf{D}}$ obtained from a particular viewpoint do not adequately represent the complete dynamic information in the 4D space, we design temporal-spatial similarity fusion module (TSSF) to adaptively select similar dynamic information from decoupled features under different views at the same timestamp.
@@ 3.4 Temporal-Spatial Similarity Fusion (TSSF)
Inspired by [49], we retrieve the decoupled features from Gaussian points by view projection. Then we integrate decoupled features into point features based on the given camera pose. The Gaussian points are well-aligned with decoupled features thanks to the projection operations. Therefore, point features $\textbf{F}_{\textbf{p}}=\{f_{p}^{(0,0)},f_{p}^{(0,1)},\dots,f_{p}^{(i,j)},\dots,f_{p}^{(t,v)}\}$ encapsulate extensive dynamic information from decoupled features. In fact, among different viewpoints, the texture, shape, and motion representations from the same spatial area of the object are similar in the same timestamps. Based on this fact, utilizing the semantic similarities in point features across viewpoints and aggregating them contributes to the enhancement of dynamic representations in point features. Hence, we propose an adaptive fusion in Fig.14.
@@
At the same timestamp, we design global awareness fusion (GA) in adaptive fusion to merge similar dynamic information from point features along spatial axes. However, under conditions of significant differences among multi-views, the effect of merging in GA is affected by that dynamic regions are noticeably obscured in specific viewpoints. Therefore, we need to reduce the impact of point features under these views for merging while retaining the useful dynamic information in these features. Based on this condition, we design distance awareness fusion (DA).
Global awareness (GA). At time $i$, the fully-connection layer (FC) generates score maps $\textbf{W}=\{w^{(0,0)},\dots,w^{(t,v)}\}$ for the point features of each view,
@@
Distance awareness (DA). Frames from single-view video are regarded as front view. The front frame reveals real and visible motion areas of the object at various timestamps compared to the other views. This information within the front frame requires preservation as much as possible while supplementing the useful dynamic information from other views for accurancy motion prediction. Thus, at time $i$, we calculate the $L_{1}$ distance between point features from the front view and point features from the other views. The distance quantifies the difference information between point features extracted from the front view and those derived from other viewpoints. The difference information contains dynamic information that cannot be represented by the point features of the front view. Then, we append point features from other views with distance features and merge them by Eq. 3 and Eq. 4. The fused point features from other views are appended to point features from the front view. We merge these features by Eq. 3 and Eq. 4 to acquire the fused point features $f_{a}^{i}$. Through the above pre-processing, we reduce the impact of point features under other views according to the front view while retaining useful dynamic information in point features under other views.
In addition to fused features through adaptive fusion, we also present dynamic Gaussian features as 3D Gaussian features with HexPlane [50] for inherently regularizing features of fields and guaranteeing their smoothness. Formally, the dynamic Gaussian features at time $i$ can be defined as:
@@
where $X_{i}=(x_{i},y_{i},z_{i})$ denotes the position of Gaussian points, $s_{i}$ and $\gamma_{i}$ denote scale and rotation at time $i$. $\sigma$ and $\zeta$ represent opacity and spherical harmonic coefficients of the radiance. Then, we combine dynamic Gaussian features $f_{\text{hg}}^{i}$ and fused point features $f_{a}^{i}$ at time $i$, and map them to fused Gaussian features via learnable linear transformation. The fused Gaussian features represent Gaussian intrinsic properties while capturing rich dynamic information. Therefore, fused Gaussian features are essential for accurately predicting deformation using Deformation MLP.